A man faces north and covers 7km, turns West and covers 2km then turns South and covers 4km, then turns West again and covers 2km. He is ______ from his starting point.
Answer by rfer(16322)

he is now 4 west an 3 north
a^2+b^2=c^2
4^2+3^2=c^2
16+9=25
sqrt25=5 km from start